Suppose an investment offers to quadruple your money in 36 months (don’t believe it). What rate of return per quarter are you being offered? (Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text ta copy ▼ в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 C Conditional Format CeInsert Delete Format Formatting, as Table w styles. ▼ ㆆ ▼ Sort &Find & 2 ClearFe Select Edting Format Painter Clipboard Alignment Number Cells NU NV NW NX NY NZ OA OB OC OD OE OF OG 4 5 6 PV= FV= MONTHS YEARS RATE 100 36 58.74% excel-RATE(Nw7,,-NVV4,NWS) 9 10 rate of return per quarter 12 13 14 15 16 17 18 19 20 21 (1 +r/m)m -1 EAR- 58.74%- (1+r/4)^4-1 1.5874 (1+r/4)A4 1.587441/4)-(1+r/4) 1.1224619 = 1+r/4 0.1224619 -r/4 0.1224619 r/4= 12.25% RATE OF RETURN PER QUARTER BOX DILUTION DOLLAR CoST AVGMORTGAGE EXPO Sheeti Sheet2 Shel 福 130% 제 トト11 CALCULATOR | LOAN OPTIONS fv. ann CAP STRU VALUE